# Multimodal LLM Guided Exploration and Active Mapping using Fisher Information

### v2

@@ 3.3 Information-driven Path Proposal
as a preliminary selection metric to form our final candidate target poses set $\mathcal{T}_{F}$. Thus, we can identify multiple coarse directions for explorations where we can propose multiple paths for detailed path planning and path selection. FisherRF [31] uses ${\mathbf{H}}^{\prime\prime}[{\bf w}|D^{train}]$ as an approximation for the observed Fisher Information $\mathcal{I}({\bf w})$ by computing the Hessians on the training set. This is also known as empirical Fisher Information, whose limitations have been widely discussed by Kunstner et al. [35] and Marten et al. [43]. In most scenarios, this is a reluctant design choice because the distribution of ${\bf x}\sim p({\bf x})$ is unknown (i.e., the distribution of all possible images). However, ${\bf x}\sim p({\bf x})$ in our case is tractable because it represents the possible locations where we can take an observation for the environment, a.k.a. the free space of our map. Therefore, unlike FisherRF [31] as described in Eq. 3, we propose to use Monte-Carlo sampling to compute the Fisher Information of the current model
@@
where ${\bf x}_{k}$ is drawn from a uniform distribution of camera poses in the free space of the current map which we use to approximate $p$. Besides, we also uniformly initialize 3D Gaussians in the space, which will be subsequently updated with rendering losses for visited regions. In Fig. 2, we show the relationship between PSNR and EIG on sampled poses in the Cantwell scene. The result aligns with our expectation that EIG should decrease as PSNR increases. We also show some cases on the lower-left and upper-right of the scatter plot. For the lower left capture, even though the left part is poorly reconstructed, most of the view is occupied by a textureless wall, leading to a low EIG. For the upper right capture, even though the scene has a moderate reconstruction, the content is rich, so our algorithm returns a high EIG.
Finally, we compute paths towards each pose in $\mathcal{T}_{F}$ with the A* algorithm [24] using the occupancy map, selecting which path to follow as described in Section 3.4. The path can be defined as an ordered set of camera poses from the current location $x_{t}$ at exploration step $t$ to the frontier points $x_{T}^{j}$ from the long horizon planning.
@@
UPEN [22] FBE [82] Ours
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative Comparison for Final Scene Reconstruction on Habitat-Matterport 3D Dataset mscxX4KEBcB (top), oPj9qMxrDEa (middle) and QKGMrurUVbk (bottom) scenes. We provide top-down rendering for different methods.
3.4 Short-term Path Selection with Localization Uncertainty
The EIG for 3D Gaussian parameters along a path $\mathcal{P}_{j}$ can be computed as the sum over the path of the following term [31]:
$\mathcal{I}_{\mathcal{P}_{j},i}({\bf w})={\mathbf{H}}^{\prime\prime}[{\bf w}]+\sum_{{\bf x}_{t}\in\mathcal{P}_{j},t<i}{\mathbf{H}}^{\prime\prime}[{\bf w}|{\bf x}_{t}].$ (9)
@@ 3.4 Short-term Path Selection with Localization Uncertainty
where ${\mathbf{H}}^{\prime\prime}[{\bf x}]$ is short for $-\nabla^{2}\log p(w)$ for clarity. If solely maximizing the $\mathtt{EIG}$, the robot will be more likely to explore unvisited regions. However, exploring regions that have not been well reconstructed also means the agent would have the risk of worse localization accuracy due to noise and ambiguities in the unreconstructed regions during pose optimization. The cost of localization must thus be considered during path planning to balance the importance of exploring new environments with maintaining localization accuracy. We propose to use Fisher Information as a measurement for the localization uncertainty that is also necessary for effective path planning for active Mapping algorithms. During optimization, we essentially optimize on the logarithmic mapping of $\tau_{i}\triangleq\text{Log}({\bf x}_{i})$ of our camera pose. By the Cramér–Rao bound, the covariance of $\tau_{i}\in\mathfrak{se}(3)$ can be lower-bounded with the inverse of Fisher Information matrix $\mathcal{I}(\tau_{i})$:
$\mathrm{Cov}(T(\hat{\tau_{i}}))\geq\mathcal{I}(\tau_{i})^{-1}$ (10)
where $T(\tau_{i})$ is an unbiased estimator for $\tau$ solved by iteratively optimizing photo-metric loss. Hence, we can define the localization cost $C_{\textit{local}}$ at a pose ${\bf x}_{i}$ in terms of $\tau_{i}$ as:
$C_{\textit{local}}(\tau_{i})=\log\det(\nabla_{\tau_{i}}f(\tau_{i},{\bf w})^{T}\nabla_{\tau_{i}}f(\tau_{i},{\bf w}))$ (11)
Matsuki et al. [44] computed the Jacobians of the camera pose with respect to the mean and covariances of each gaussian $\frac{\partial{\boldsymbol{\mu}_{I}}}{\partial{{\bf x}}}$ and $\frac{\partial{\boldsymbol{\Sigma}_{I}}}{\partial{{\bf x}}}$. However, we need to compute the Jacobian of $\tau_{i}$ with respect to the rendering output:
$\nabla_{\tau_{i}}f(\tau_{i},{\bf w})=\frac{\partial{f(\tau_{i},{\bf w})}}{\partial{\tau_{i}}}=\begin{bmatrix}\frac{\partial{f(\tau_{i},{\bf w})}}{\partial{\boldsymbol{\mu}_{I}}}&\frac{\partial{f(\tau_{i},{\bf w})}}{\partial{\boldsymbol{\Sigma}_{I}}}\end{bmatrix}\begin{bmatrix}\frac{\mathcal{D}{\boldsymbol{\mu}_{C}}}{\mathcal{D}{\tau_{i}}}\\ \frac{\mathcal{D}{\mathbf{W}}}{\mathcal{D}{\tau_{i}}}\end{bmatrix}$ (12)
Unlike 3DGS parameters, our computation of the localization cost can be directly computed on the full Fisher Information matrix without using the Laplace approximation [12, 42]. That is because the full Jacobian $\nabla_{\tau_{i}}f(\tau_{i},{\bf w})$ is linear to the number of Gaussians and can be efficiently computed with our CUDA implementation.
Without loss of generality, the path of exploration can be selected by minimizing the total cost for all viewpoints ${\bf x}_{i}$ along a path $\mathcal{P}_{j}$:
where $\eta$ is a hyper-parameter controlling the importance between EIG and localization accuracy. The agent can then explore the environment with planned path $\mathcal{P}$. Our active mapping system constantly updates the map, and we replan using our active path planning algorithm if we detect the agent is getting close to a possible obstacle or upon reaching the end of the previously selected path.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth MAE $\downarrow$ RMSE ATE $\downarrow$ Coverage ($m^{2}$) $\uparrow$
ANS 16.34 0.6818 0.3923 0.3886 0.1105 10.49
Active-INR (gt) 22.66 0.7652 0.2164 0.1528 - 9.20
UPEN (gt) 21.31 0.7325 0.2714 0.1696 - 8.79
UPEN 16.44 0.6678 0.4134 0.4841 0.5158 8.58
ExplORB 18.99 0.7175 0.3994 0.2664 0.2296 9.00
FBE 21.45 0.7618 0.2126 0.1028 0.1680 10.64
Ours 23.28 0.8067 0.2507 0.0696 0.0226 10.88
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth MAE $\downarrow$ RMSE ATE $\downarrow$ Coverage ($m^{2}$) $\uparrow$
UPEN (gt) 15.58 0.5175 0.3936 0.4548 - 12.69
UPEN 12.23 0.4795 0.5157 0.7356 0.4393 10.72
FBE 15.80 0.5952 0.4392 0.4085 1.2004 15.69
Ours 19.86 0.7127 0.4122 0.1666 0.0336 18.48
Refer to caption Refer to caption
Ground truth Estimated
Figure 6: Qualitative Comparison for Trajectories on Cantwell from Gibson Dataset. Left is the ground truth trajectory, right is the estimated trajectory.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth MAE $\downarrow$ RMSE ATE $\downarrow$ Coverage ($m^{2}$) $\uparrow$
w.o. LLM & Localization Uncertainty 16.15 0.6550 0.6193 0.3409 0.2478 16.06
w.o. LLM 16.94 0.6799 0.5847 0.2887 0.1694 17.26
Ours 19.86 0.7127 0.4122 0.1666 0.0336 18.48
Table 3: Ablation Study of Localization Uncertainty Term on Scenes from the HM3D Dataset. We compare our method with and without the localization uncertainty term to validate that including it provides improvements on both localization and reconstruction
4 Experiments
@@ 4.1 Experimental Set-up
Following previous methods [22, 85], our algorithm is evaluated in the Habitat Simulator [73] on the Gibson [81] and Habitat-Matterport 3D (HM3D) [56] datasets comprised of indoor scenes reconstructed from scans of real houses. We use all the scenes in the val split for the Gibson dataset. We also evaluate our method on single-floor scenes whose scene ID can be found in the supplementary. We adopt the default start point in the Habitat Simulator as the starting point for exploration in each scene. The total number of steps for each experiment is 2000. The system takes color and depth images at the resolution of 800x800 and outputs a discrete action at each step. The action space consists of MOVE FORWARD by 5cm, TURN LEFT, and TURN RIGHT by 5°$\degree$. The field of view (FOV) is set to 90°$\degree$ vertically and horizontally. Please refer to the supplement for more details about the evaluation split and other hyper-parameters.
@@
We evaluated our method using the Peak-signal-to-noise ratio (PSNR), Structural Similarity Index Measure  (SSIM) [79], Learned Perceptual Image Patch Similarity (LPIPS) [89] for RGB rendering and mean absolute error (MAE) for depth rendering as metrics for scene reconstruction quality. We calculate these metrics using 2000 points uniformly sampled from the movement plane of the agent in the scene, discarding any points that are not navigable. We argue that the rendering quality reflects both reconstruction quality and pose accuracy because high tracking accuracy would help the training of the 3D Gaussian Splatting model. Meanwhile, misaligned poses will lead to misaligned rendering at test time thus leading to inferior results. Following previous approaches [22, 7], we also use coverage in $m^{2}$ as an evaluation metric. To evaluate the pose estimation accuracy, we use the root mean squared average tracking error (RMSE ATE), but as the trajectories for each method are different, the RMSE ATE should only be considered along with other metrics such as coverage.
@@
We compare to two exploration methods which assume ground truth pose: UPEN [22] and Active Neural Mapping (active-INR) [85]. UPEN chooses trajectories to reduce the epistemic uncertainty of the occupancy prediction produced by an ensemble of models. Active-INR aims to minimize the prediction robustness against random weight perturbation of its signed distance field scene representation. We also compare our method with Active Neural SLAM (ANS) [7], explORB [53] and Frontier Based Exploration (FBE) [82] without ground truth pose provided. ANS [7] learns to predict a map and estimate a global goal based on it. ExplORB [53] computes the Fisher Information of the Hessian on its pose graph optimization.
To compare the rendering quality fairly, we run all the baselines using the MonoGS [44] backend for reconstruction. We run UPEN and FBE online, but for ANS, active-INR, and ExplORB, we record and playback trajectories obtained using their source code. Because the forward step size for ANS is much larger than for our method, we interpolate the trajectory so that the forward step size matches our method’s to make the steps comparable. For ExplORB, since the official implementation is based on MoveBase, which uses velocity commands, we sample the trajectory at 5 Hz. We also found that ANS, active-INR, and UPEN failed on some scenes due to localization failure. ANS produces a pose estimate (using information from noisy pose sensors not provided to our pipeline), so we set the pose estimate of the MonoGS backend to the one from ANS. As active-INR and UPEN do not produce a pose estimate, we evaluate them using the ground-truth pose.
@@ 4.2 Comparison Against Previous Methods
Table 1 shows the results of our method and the baselines for exploration in scenes from the Gibson dataset [81], and Table 2 shows the results of our method and some baselines on HM3D [6]. Our method outperforms the baselines on all metrics. We further qualitatively compare the reconstruction qualities after active exploration in Fig. 4 and Fig. 5, and the trajectories in Fig. 6. Our method does not have major errors from failed localization, and we have fewer gaps in the scenes than other methods. For example, in the Ribera scene, all methods except for us and FBE miss the bathroom at the bottom left, and FBE misses more of the area around the sofa than us. For the trajectories, we show the estimated and ground truth trajectories for the Cantwell scene from the Gibson dataset. Cantwell is a relatively large and challenging scene, so it is suitable for showing the differences between methods. We show only a few baselines to keep the figure legible. We can see that Active-INR stays in a smaller area than the rest of the methods. ANS often goes close to walls, whereas FBE and our method are generally more toward the center of corridors or rooms. ANS also extends further into rooms than these methods. FBE does not go to the region on the bottom right, and ANS does not go to the region on the bottom left, whereas our method efficiently visits most areas of the scene.
4.3 Ablation Study
To validate the effectiveness of localization uncertainty and LLM, we performed an ablative study of each component in our system on the Gibson dataset in Table. 3. As can be seen, the average trajectory error is much lower with the localization uncertainty than without. In addition, after adding LLM, the model performs better in the overall rendering. The ATE slightly increases after adding LLM, but this could be attributed to the larger area traversed by the robot.
@@ 5 Conclusion
We present an active mapping system that can autonomously explore an environment, the first method to do so using a 3D Gaussian representation without ground truth camera poses. By rendering maps using our scene representation, we can seamlessly query a multimodal Large Language Model for long-horizon planning to leverage its knowledge of scene layouts. This allows the robot to efficiently explore the scene while avoiding setting unreachable navigation goals. We then plan an optimal path to the long-horizon goal using our information-driven path proposal and selection algorithm, which balances the information gain with respect to the map against the localization uncertainty, ensuring that the robot captures the geometric details of the scene for a high-quality reconstruction. We evaluate our method on scenes from the Gibson [81] and Habitat-Matterport 3D  [56] datasets in terms of the rendering quality, coverage, and average tracking error. We compare to recent exploration methods with [22, 85] and without [53, 7] ground-truth pose, using the same scene representation for reconstruction as us for fairness, and show that our method has superior performance. To enable our method to support more robotics applications, future work could extend it to take advantage of our 3D scene representation and consider movement with higher degrees of freedom (DOF) than the currently supported 3DOF. Incorporating semantic features [91, 68, 55] to allow for grounding language to the scene would also enable many robotics and computer vision applications.

### v3

@@ 3.3 Expected Information Gain for Path Proposal and Selection
as a preliminary selection metric to form our final candidate target poses set $\mathcal{T}_{F}$. We calculate $\mathcal{I}({\bf w})$ differently to FisherRF [33], which approximates it using ${\mathbf{H}}^{\prime\prime}[{\bf w}|D^{train}]$ by computing the Hessians on the training set. This is also known as empirical Fisher Information, whose limitations have been widely discussed [40, 48]. In most scenarios, this is a reluctant design choice because the distribution of ${\bf x}\sim p({\bf x})$ is unknown (i.e., the distribution of all possible images). However, ${\bf x}\sim p({\bf x})$ in our case is tractable because it represents the possible locations where we can take an observation for the environment, i.e. the free space of our map. Therefore, unlike FisherRF [33] as described in Eq. 3, we propose to use Monte-Carlo sampling to compute the Fisher Information of the current model
@@
where ${\bf x}_{k}$ is drawn from a uniform distribution of camera poses in the free space of the current map which we use to approximate $p$. We also uniformly initialize 3D Gaussians in the space, which will be subsequently updated with rendering losses for visited regions.
After the set of final candidate poses $\mathcal{T}_{F}$ has been formed, we construct paths to each candidate pose with the A* algorithm [25] using the occupancy map. The path can be defined as an ordered set of camera poses from the current location $x_{t}$ at exploration step $t$ to the frontier points $x_{T}^{j}$ from the long horizon planning.
@@
To evaluate how beneficial a path $\mathcal{P}_{j}$ would be for improving the reconstruction of the scene we consider the EIG for the 3D Gaussian parameters along that path, which can be computed as the sum over the path of the following term [33]:
$\mathcal{I}_{\mathcal{P}_{j},i}({\bf w})=\mathcal{I}({\bf w})+\sum_{{\bf x}_{t}\in\mathcal{P}_{j},t<i}\mathcal{I}({\bf w}|{\bf x}_{t}).$ (9)
@@
UPEN [23] FBE [89] Ours UPEN [23] FBE [89] Ours
Figure 5: Qualitative Comparison for Final Scene Reconstruction on Habitat-Matterport 3D Dataset oPj9qMxrDEa (left) and QKGMrurUVbk (right) scenes. We provide top-down rendering for different methods.
3.4 Localization Uncertainty for Path Selection
If solely maximizing the EIG, the robot will be more likely to explore unvisited regions. However, exploring regions that have not been well reconstructed also means the agent would have the risk of worse localization accuracy due to noise and ambiguities in the unreconstructed regions during pose optimization. The cost of localization must thus be considered during path planning to balance the importance of exploring new environments with maintaining localization accuracy. We propose to use Fisher Information as a measurement for localization uncertainty to address this. During optimization, we essentially optimize on the logarithmic mapping $\tau_{i}\triangleq\text{Log}({\bf x}_{i})$ of our camera pose. By the Cramér–Rao bound, the covariance of $\tau_{i}\in\mathfrak{se}(3)$ can be lower-bounded with the inverse of Fisher Information matrix $\mathcal{I}(\tau_{i})$:
$\mathrm{Cov}(T(\hat{\tau_{i}}))\geq\mathcal{I}(\tau_{i})^{-1}$ (10)
where $T(\tau_{i})$ is an unbiased estimator for $\tau_{i}$ solved by iteratively optimizing photo-metric loss. Hence, we can define the localization cost $C_{\textit{local}}$ at a pose ${\bf x}_{i}$ in terms of $\tau_{i}$ as:
$C_{\textit{local}}(\tau_{i})=-\log\det(\nabla_{\tau_{i}}f(\tau_{i},{\bf w})^{T}\nabla_{\tau_{i}}f(\tau_{i},{\bf w}))$ (11)
Matsuki et al[49] computed the Jacobians of the camera pose with respect to the mean and covariances of each gaussian $\frac{\partial{\bm{\mu}_{I}}}{\partial{{\bf x}}}$ and $\frac{\partial{\bm{\Sigma}_{I}}}{\partial{{\bf x}}}$. However, we need to compute the Jacobian of $\tau_{i}$ with respect to the rendering output:
$\nabla_{\tau_{i}}f(\tau_{i},{\bf w})=\frac{\partial{f(\tau_{i},{\bf w})}}{\partial{\tau_{i}}}=\begin{bmatrix}\frac{\partial{f(\tau_{i},{\bf w})}}{\partial{\bm{\mu}_{I}}}&\frac{\partial{f(\tau_{i},{\bf w})}}{\partial{\bm{\Sigma}_{I}}}\end{bmatrix}\begin{bmatrix}\frac{\mathcal{D}{\bm{\mu}_{C}}}{\mathcal{D}{\tau_{i}}}\\ \frac{\mathcal{D}{\mathbf{W}}}{\mathcal{D}{\tau_{i}}}\end{bmatrix}$ (12)
Unlike 3DGS parameters, our computation of the localization cost can be directly computed on the full Fisher Information matrix without using the Laplace approximation [14, 47]. That is because the full Jacobian $\nabla_{\tau_{i}}f(\tau_{i},{\bf w})$ is linear to the number of Gaussians.
We select the best path by minimizing the total cost for all viewpoints ${\bf x}_{i}$ along a path $\mathcal{P}_{j}$:
where $\eta$ is a hyper-parameter controlling the relative importance between EIG and localization accuracy. Our active mapping system constantly updates the map, and we replan using our active path planning algorithm if we detect the agent is getting close to a possible obstacle or upon reaching the end of the previously selected path.
4 Experiments
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth MAE $\downarrow$ RMSE ATE $\downarrow$ Completeness (%) $\uparrow$
ANS [9] 16.34 0.6818 0.3923 0.3886 0.1105 35.33
UPEN [23] 16.44 0.6678 0.4134 0.4841 0.5158 22.66
ExplORB [58] 18.99 0.7175 0.3994 0.2664 0.2296 30.23
FBE [89] 21.45 0.7618 0.2126 0.1028 0.1680 55.87
Ours 23.28 0.8067 0.2507 0.0696 0.0226 84.38
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth MAE $\downarrow$ RMSE ATE $\downarrow$ Completeness (%) $\uparrow$
UPEN [23] 12.23 0.4795 0.5157 0.7356 0.4393 17.48
ExplORB (gt) [58] 17.81 0.3694 0.6810 0.5071 - 31.92
FBE [89] 15.80 0.5952 0.4392 0.4085 1.2004 22.42
Ours 19.86 0.7127 0.4122 0.1666 0.0336 49.76
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth MAE $\downarrow$ RMSE ATE $\downarrow$ Completeness (%) $\uparrow$
w.o. LLM & Localization Uncertainty 16.15 0.6550 0.6193 0.3409 0.2478 35.40
w.o. LLM 16.94 0.6799 0.5847 0.2887 0.1694 37.26
Ours (Llava-7b) 18.46 0.6805 0.4623 0.2033 0.0159 17.41
$\lambda=2\times 10^{-6}$ $\dagger$ 18.90 0.6976 0.4408 0.1966 0.0479 18.12
$\lambda=5\times 10^{-6}$ $\dagger$ 18.90 0.6946 0.4602 0.2145 0.0551 18.05
@@ 4 Experiments
Table 3: Ablation Study of Localization Uncertainty Term on Scenes from the HM3D Dataset. We compare our method with and without the localization uncertainty term to validate that including it improves localization and reconstruction. $\dagger$: the multimodal LLM is GPT-4o in the experiment.
@@ 4.1 Experimental Set-up
[…]
@@
We evaluated our method using the Peak-signal-to-noise ratio (PSNR), Structural Similarity Index Measure  (SSIM) [86], Learned Perceptual Image Patch Similarity (LPIPS) [96] for RGB rendering and mean absolute error (MAE) for depth rendering as metrics for scene reconstruction quality. We calculate these metrics using 2000 points uniformly sampled from the movement plane of the agent in the scene, discarding any points that are not navigable. We argue that the rendering quality reflects both reconstruction quality and pose accuracy because high tracking accuracy would help the training of the 3D Gaussian Splatting model. Meanwhile, misaligned poses will lead to misaligned rendering at test time thus leading to inferior results. Following previous approaches [92, 39], we also use the completion ratio as an evaluation metric. To evaluate the pose estimation accuracy, we use the root mean squared average tracking error (RMSE ATE), but as the trajectories for each method are different, the RMSE ATE should only be considered along with other metrics such as completness.
@@
We compare to two exploration methods which assume ground truth pose: UPEN [23] and Active Neural Mapping (active-INR) [92]. We also compare our method with Active Neural SLAM (ANS) [9], explORB [58] and Frontier Based Exploration (FBE) [89] without ground truth pose provided. Note that we do not run ANS on HM3D as it is not trained on this dataset.
To compare the rendering quality fairly, we run all the baselines using the MonoGS [49] backend for reconstruction. We run UPEN and FBE online, but for ANS, active-INR, and ExplORB, we record and playback trajectories obtained using their source code. Because the forward step size for ANS is much larger than for our method, we interpolate the trajectory so that the forward step size matches our method’s to make the steps comparable. For ExplORB, since the official implementation is based on MoveBase, which uses velocity commands, we sample the trajectory at 5 Hz. We also found that ANS, active-INR, and UPEN failed on some scenes due to localization failure of the MonoGS backend. ANS produces a pose estimate (using information from noisy pose sensors not provided to our pipeline), so we set the pose estimate of the MonoGS backend to the one from ANS. As active-INR and UPEN do not produce a pose estimate, we evaluate them using the ground-truth pose.
Figure 6: Runtime Analysis of the Active Mapping System. We provide running time statistics by each step and by the percentage of total time consumed on one episode.
Figure 7: Qualitative Comparison for Trajectories on Cantwell from Gibson Dataset.
@@ 4.2 Comparison Against Previous Methods
Table 1 shows the results of our method and the baselines for exploration in scenes from the Gibson dataset [88], and Table 2 shows the results on HM3D [8]. Our method outperforms the baselines on all metrics. We further qualitatively compare the reconstruction qualities after active exploration in Fig. 4 and Fig. 5, and the trajectories in Fig. 7. Our method does not have major errors from failed localization, and we have fewer gaps in the scenes than other methods. For example, in the Ribera scene, all methods except for us and FBE miss the bathroom at the bottom left, and FBE misses more of the area around the sofa than us. We show the trajectories taken for the Cantwell scene from the Gibson dataset, which is a relatively large and challenging scene. We show only a few baselines to keep the figure legible. We can see that ANS [9] does not go to the region on the bottom left, Active-INR [92] stays in the center region, and FBE does not go to the region on the bottom right, whereas our method efficiently visits most areas of the scene.
4.3 Ablation Study and Runtime Analysis
To validate the effectiveness of localization uncertainty, hyperparameters, and LLM, we performed an ablative study of each component in our system on the HM3D dataset in Table. 3. As can be seen, the average trajectory error is much lower with the localization uncertainty than without. In addition, after adding LLM, the model performs better in all metrics. We also noticed the tracking error was lower after introducing multimodal LLM to the active mapping system. We hypothesize this improvement is because improved scene reconstruction can also help reduce tracking errors.
We additionally provide a breakdown of the runtime of our method, shown in Fig.˜6.
@@ 5 Conclusion
We present an active mapping system that can autonomously explore an environment, the first method to do so using a 3D Gaussian representation without ground truth camera poses. By rendering maps using our scene representation, we can seamlessly query a multimodal Large Language Model for long-horizon planning to leverage its knowledge of scene layouts. This allows the robot to efficiently explore the scene while avoiding setting unreachable navigation goals. We then plan an optimal path to the long-horizon goal using our information-driven path proposal and selection algorithm, which balances the information gain with respect to the acene against the localization uncertainty, ensuring that the robot captures the geometric details of the scene for a high-quality reconstruction. We evaluate our method on scenes from the Gibson [88] and Habitat-Matterport 3D [61] datasets in terms of the rendering quality, completeness, and average tracking error. To enable our method to support more robotics applications, future work could extend it to take advantage of our 3D scene representation and consider movement with higher degrees of freedom (DOF) than the currently supported 3DOF. Our method could also be extended to task-driven exploration such as finding objects [2, 37], leveraging the LLM for efficient search. Incorporating semantic features [98, 73, 60] to allow for grounding language to the scene would also enable many robotics and computer vision applications.